Jane drove 230 miles at an average speed of 50 miles per hour. How many minutes was Jane driving?
Answer by TimothyLamb(4379)

s = d/t
t = d/s
t = 230/50
t = 4.6 hours
---
time in minutes:
(4.6 hours) * (60 minutes/hour) = 276 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php